OpenStudy (anonymous):

5. The discriminant of a quadratic equation is 5. Which describes the number and type of solutions of the equation? I think its 2 complex solutions

OpenStudy (cwrw238):

no its 2 solutions but they are both rel numbers

OpenStudy (cwrw238):

they would be complex if discriminant was negative 5

OpenStudy (anonymous):

if you mean 2 real solutions your wrong Im redoing my test and I got that to but it was wrong

OpenStudy (anonymous):

The discriminant of a quadratic equation is 5. Which describes the number and type of solutions of the equation?

OpenStudy (cwrw238):

there are 2 solutions - do you agree with that?

OpenStudy (cwrw238):

no t didnt say there are 2 real and 2 complex. the answer is definitely 2 real solutions

OpenStudy (anonymous):

Solve x^2 + x + 5 = 0 using the Quadratic Formula. What is the solution set? okay I'll put that one can you help me with this one plz

OpenStudy (cwrw238):

x = [-1 +/- sqrt(1^2 - 4*1*5)] / 2

OpenStudy (cwrw238):

note the discriminant here : - its 1 - 20 = -19 - here we have 2 complex number roots

OpenStudy (anonymous):

ok I got 1\[\left\{ 1\pm i/2 \right\}\]

OpenStudy (cwrw238):

x = -1 +i sqrt19 -1 - i sqrt19 ---------- , ------------ 2 2

OpenStudy (anonymous):

how did you got sqrt of 19? Im confuse

OpenStudy (cwrw238):

the discriminant is b^2 - 4ac for the equation ax^2 + bx + c = 0 compare this with your equation : b = 1, a = 1 and c = 5 so b^2 - 4ac = 1^2 - 4*1*5 = 1 - 20 = -19

OpenStudy (anonymous):

on the first question it was 2 complex solutions for -5 you were not correct and on the second one you were correct and thank you!

OpenStudy (cwrw238):

sqrt (-19) = sqrt( -1 * 19) = i sqrt19

OpenStudy (cwrw238):

yw